OpenStudy (anonymous):

Solve the equation by completing the square. If necessary, round to the nearest hundredth. x² – 18x = 19 1; 19 –1; 19 3; 6 –3; 1 Could someone help me with this? Thanks in advance.

OpenStudy (anonymous):

What does that last pic say?

OpenStudy (anonymous):

who needs ... square?!

OpenStudy (anonymous):

thanks though :D

OpenStudy (anonymous):

What do you think @Dharmaputra_993?

OpenStudy (anonymous):

I was typing the answer, but Luis is helping you so it's okay :)

OpenStudy (anonymous):

Okay, Thanks for trying ;D

OpenStudy (anonymous):

@Luis_Rivera Wouldnt it be \[(x + 1)(x - 19) = 0\]

OpenStudy (anonymous):

Can you explain you reasoning?

jimthompson5910 (jim_thompson5910):

x^2 - 18x = 19 x^2 - 18x - 19 = 0 Find two numbers that multiply to -19 and add to -18...those two numbers are 1 and -19, so x^2 - 18x - 19 factors to (x+1)(x-19) x^2 - 18x - 19 = 0 (x+1)(x-19) = 0 x+1 = 0 or x-19 = 0 ... use the zero product property here x = -1 or x = 19

OpenStudy (anonymous):

Can you explain your* reasoning?

OpenStudy (anonymous):

@some_someone ?

OpenStudy (anonymous):

\[x² – 18x = 19\] \[x² – 18x + 81= 100\] \[(x - 9)^2 = 100\] \[x - 9 = 10 \] \[x = 19\] \[x - 9 = -10 \] \[x = -1\]

OpenStudy (anonymous):

Well, it can't be 19...

OpenStudy (anonymous):

Oops...

OpenStudy (anonymous):

Lol brain fart :)P

OpenStudy (anonymous):

@jim_thompson5910 so i was right? and @FlashMobRiot it would be –1; 19

OpenStudy (anonymous):

Sorr, I'm multitasking :)

OpenStudy (anonymous):

its okaii :)

OpenStudy (anonymous):

Yup, you and @Luis_Rivera and @jim_thompson5910 were all right :D

jimthompson5910 (jim_thompson5910):

yeah the solutions to x^2 - 18x = 19 are x = -1 or x = 19

jimthompson5910 (jim_thompson5910):

Luis_Rivera was close, but he had his signs backward

OpenStudy (anonymous):

I can't give you all best response...so try to be as unbiased as you can, who should get it?

OpenStudy (anonymous):

I think @some_someone maybe should get it because he has the lowest number compared to @jim_thompson5910

OpenStudy (anonymous):

Is that ok? @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

go for it, i don't mind

OpenStudy (anonymous):

Ok thanks :)P

jimthompson5910 (jim_thompson5910):

as long as i helped, that's all i care about really

OpenStudy (anonymous):

Yup you definetly did

jimthompson5910 (jim_thompson5910):

I'm glad I did

OpenStudy (anonymous):

I have one other question I'm a bit skeptical about, do you mind checking it out? If so here it is: Solve. x² – 81 = 0 1) 0 2) –9 *3) –9, 9 4) 9 (Hint: * means the one I chose)

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

if you're not sure, just plug in the possible solutions back in for x

OpenStudy (anonymous):

Ok thanks for looking it over

OpenStudy (anonymous):

Oh Ok

jimthompson5910 (jim_thompson5910):

then evaluate and simplify and you should get 0 = 0, which is a true equation

OpenStudy (anonymous):

Yea...I do that alot on my state test lol. Well thanks have a great day, and bye :D

OpenStudy (anonymous):

Once @some_someone is done typing I'll close the question.

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

that's good to know checking your answer is always a good idea yw, cya later

OpenStudy (anonymous):

Thanks, Bye everyone I'm logging out.